# Easi3R: Estimating Disentangled Motion from DUSt3R Without Training

### v2

@@ 2 Related Work
SfM and SLAM. Structure-from-Motion (SfM) [41, 42, 52, 2, 48, 51] and Simultaneous Localization and Mapping (SLAM) [34, 32, 9, 13] have long been the foundation for 3D structure and camera pose estimation. These methods are done by associating 2D correspondences [5, 32, 28, 10, 47] or minimizing photometric errors [13, 12], followed by bundle adjustment (BA) [3, 59, 57, 55, 62, 6] to refine structure and motion estimates. Although highly effective with dense input, these approaches often struggle with limited camera parallax or ill-posed conditions, leading to performance degeneracy. To overcome these limitations, DUSt3R [64] introduced a learning-based approach that directly predicts two pointmaps from an image pair in the coordinate space of the first view. This approach inherently matches image features and rigid body view transformation. By leveraging a Transformer-based architecture [11] and direct point supervision on large-scale 3D datasets, DUSt3R establishes a robust Multi-View Stereo (MVS) foundation model. However, DUSt3R and the follow-up methods [61, 56, 27, 33] assume primarily static scenes, which can lead to significant performance degradation when dealing with videos with dynamic objects.
@@ 3.1 DUSt3R with Dynamic Video
DUSt3R is designed for pose-free reconstruction, taking two RGB images - $I^{a},I^{b}$, where $a,b\in[1,\dots,T]$ as input and output two pointmaps in the reference view coordinate space, $X^{a\rightarrow a},X^{b\rightarrow a}\in\mathbb{R}^{W\times H\times 3}$:
@@ 3.2 Secrets Behind DUSt3R
Temporal attention maps. In the following, we extend the above single-pair formulation to multiple pairs to explore their temporal attention correlations. For a specific frame $I^{t}$, it pairs with multiple frames, resulting in $2(n-1)$ attention maps per frame. As shown in the upper row of Figure 2, window size of 3 corresponds to 4 pairs. To aggregate the pairwise cross-attention maps temporally, we compute the mean and variance over pairs that the view serves as source and reference:
@@ 3.3 Dynamic Object Segmentation
we then obtain per-frame dynamic object segmentation $\mathbf{M}^{t}=\left[\mathbf{A}^{t=\text{dyn}}>\alpha\right]$ using  Eq. 9 and $\mathbf{M}^{t}\in\mathbb{R}^{h\times w}$, $\alpha$ denotes a pre-defined attention threshold and the []delimited-[][$\cdot$][ ⋅ ] is the Iverson bracket. Note that the segmentation is processed frame by frame. To enhance temporal consistency, we apply a feature clustering method that fuses information across all frames; see the supplementary materials for more details.
@@
$\mathrm{softmax}(\tilde{\mathbf{A}}_{l}^{a\leftarrow b})=\begin{cases}0&\text{if }\mathbf{M}^{a\leftarrow b}\\ \mathrm{softmax}(\mathbf{A}_{l}^{a\leftarrow b})&\text{otherwise}\end{cases}$ (10)
@@
Global alignment. We align the predicted pointmaps from the sliding windows with the global world coordinate using Eq. 2. Moreover, thanks to dynamic region segmentation, our method also supports segmentation-aware global alignment with optical flow. In particular, we incorporate a reprojection loss to ensure that the projected point flow remains consistent with the optical flow estimation [66]. Specifically, given an image pair $(a,b)$, we compute the camera motion from frame $a$ to frame $b$, denoted by ^$\hat{\mathcal{F}}^{a\rightarrow b}$, by projecting the global point map $\mathcal{X}^{b}$ from camera $(\mathbf{P}^{a},\mathbf{K}^{a})$ to camera $(\mathbf{P}^{b},\mathbf{K}^{b})$. We then enforce the consistency between the computed flow and the estimated optical flow $\mathcal{F}^{a\rightarrow b}$ in static regions $(1-\mathbf{M}^{t})$:
@@ 4 Experiments
[…]
@@
[…]
@@ Appendix C Limitations
Alignment Method Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
@@
[…]

### v3

@@ 2 Related Work
SfM and SLAM. Structure-from-Motion (SfM) [41, 42, 52, 2, 48, 51] and Simultaneous Localization and Mapping (SLAM) [34, 32, 9, 13] have long been the foundation for 3D structure and camera pose estimation. These methods are done by associating 2D correspondences [5, 32, 28, 10, 47] or minimizing photometric errors [13, 12], followed by bundle adjustment (BA) [3, 59, 57, 55, 62, 6] to refine structure and motion estimates. Although highly effective with dense input, these approaches often struggle with limited camera parallax or ill-posed conditions, leading to performance degeneracy. To overcome these limitations, DUSt3R [64] introduced a learning-based approach that directly predicts two pointmaps from an image pair in the coordinate space of the first view. This approach inherently matches image features and rigid body view transformation. By leveraging a Transformer-based architecture [11] and direct point supervision on large-scale 3D datasets, DUSt3R establishes a robust Multi-View Stereo (MVS) foundation model. However, DUSt3R and the follow-up methods [61, 56, 27, 33] assume primarily static scenes, which can lead to significant performance degradation when dealing with videos with dynamic objects [76].
@@ 3.1 DUSt3Rwith Dynamic Video
DUSt3Ris designed for pose-free reconstruction, taking two RGB images - $I^{a},I^{b}$, where $a,b\in[1,\dots,T]$ - as input and output two pointmaps in the reference view coordinate space, $X^{a\rightarrow a},X^{b\rightarrow a}\in\mathbb{R}^{W\times H\times 3}$:
@@ 3.2 Secrets Behind DUSt3R
Temporal attention maps. In the following, we extend the above single-pair formulation to multiple pairs to explore their temporal attention correlations. For a specific frame $I^{t}$, it pairs with multiple frames, resulting in 2(n1)2(n-1) attention maps per frame. As shown in the upper row of Figure 2, window size of 3 corresponds to 4 pairs. To aggregate the pairwise cross-attention maps temporally, we compute the mean and variance over pairs that the view serves as source and reference:
@@ 3.3 Dynamic Object Segmentation
we then obtain per-frame dynamic object segmentation $\mathbf{M}^{t}=\left[\mathbf{A}^{t=\text{dyn}}>\alpha\right]$ using  Eq. 9 and $\mathbf{M}^{t}\in\mathbb{R}^{h\times w}$, $\alpha$ denotes a pre-defined attention threshold and the $\cdot$] is the Iverson bracket. Note that the segmentation is processed frame by frame. To enhance temporal consistency, we apply a feature clustering method that fuses information across all frames; see the supplementary materials for more details.
@@
[…]
@@
Global alignment. We align the predicted pointmaps from the sliding windows with the global world coordinate using Eq. 2. Moreover, thanks to dynamic region segmentation, our method also supports segmentation-aware global alignment with optical flow. In particular, we incorporate a reprojection loss to ensure that the projected point flow remains consistent with the optical flow estimation [66]. Specifically, given an image pair $(a,b)$, we compute the camera motion from frame $a$ to frame $b$, denoted by $\hat{\mathcal{F}}^{a\rightarrow b}$, by projecting the global point map $\mathcal{X}^{b}$ from camera $(\mathbf{P}^{a},\mathbf{K}^{a})$ to camera $(\mathbf{P}^{b},\mathbf{K}^{b})$. We then enforce the consistency between the computed flow and the estimated optical flow $\mathcal{F}^{a\rightarrow b}$ in static regions $(1-\mathbf{M}^{t})$:
@@ 4 Experiments
Figure 5: Qualitative Results of Dynamic Object Segmentation. “Ours” refers to the $\mbox{{Easi3R}}_{\text{monst3r}}$ setting. Here, we present the enhanced setting, where outputs from different methods serve as prompts and are used with SAM2 [46] for mask inference.
@@
Method Flow $uparrow JM\uparrow JR\uparrow$ $uparrow JM\uparrow JR\uparrow$ $uparrow JM\uparrow JR\uparrow$
@@ Appendix C Limitations
Alignment Method Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ $1.25\uparrow$
@@
Supervision Method RGB Optical Flow Point Tracks Mask 4D Reconstruction $\uparrow JM$-M $\uparrow$ $\uparrow JM$-M $\uparrow$